Consider a system of two particles In the xy plane
Answer #1

given that

m1 = 1.85 kg

r1 = (1i+2j )m

v1 = (3i+5j) m/s

m2 = 2.90 kg

r2 = (-4i-3j) m

v2 = (3i-2j) m/s

we know that

part(b)

center of mass of position is

Xcm = (m1*x1 + m2*x2 )/m1+m2

Xcm = (1.85*1+2.90*(-4))(1.85+2.90)

Xcm = -9.75/4.75

Xcm = -2.05 m

Ycm = (m1y1+m2y2)/m1+m2

Ycm = (1.85*2+2.90*(-3)) / (1.85+2.90)

Ycm = -5/4.75

Ycm = -1.05 m

Rcm = Xcm+Ycm

Rcm = (-2.05i-1.05j) m

part(c)

velocity of the center of mass

Vx,cm = (m1*V1x +m2*V2x)/m1+m2

Vx,cm = 1.85*3+2.90*3/1.85+2.90

Vx,cm = 3 m/s

Vy,cm = m1*V1y+m2*V2y /m1+m2

Vy,cm = 1.85*5+2.90*(-2) / ()1.85+2.90

Vy,cm = -0.72m/s

Vcm = Vx,cm+Vy,cm

Vcm = (3i -0.72 j) m/s

part(d)

total linear momentum

Px = m1*V1x+m2*V2x

Px = 1.85*3+2.90*3

Px =14.25 kg*m/s

Py = m1*V1y+m2*V2y

Py = 1.85*5+2.90*(-2)

Py = 9.25-5.8

Py =3.45 kg*m/s